OpenStudy (anonymous):

will give medal and fan if u help with 4 questions math grade 8 t over 3 minus 1 over 2 = 1

OpenStudy (usukidoll):



OpenStudy (anonymous):

?

OpenStudy (usukidoll):

It's SHeldon! :D

OpenStudy (anonymous):

sheldon?

OpenStudy (anonymous):

t/3 - 1/2 = 1 multiply both sides by 3; t - 1/2 = 3 add 1/2 to both sides; t = 4.5 does that help?

OpenStudy (anonymous):

are u sure its not 5.8?

OpenStudy (anonymous):

It should be 4.5. Could you post your work? I can try to find the error

OpenStudy (usukidoll):